# Efficient Safety Retrofitting Against Jailbreaking for LLMs

### v1

@@ 1 Introduction
A common concern across LLM domains and applications is jailbreaking [Wei et al.(2024a), Huang et al.([n. d.])]; the introduction of malicious prompts with the purpose of leading the model towards producing unsafe content to requests that, without the attack prompt, would not responded to. As of today, a wide variety of jailbreaking methods exist [Yi et al.(2024)], and more will appear as all it takes to find them is inference access to an LLM and imagination. Considering the challenges jailbreaking represents for LLM safety, this is introduced into the DPO experimentation as an fundamental aspect to study.
@@
The consistency of the above experiments is validated through an independent study on the agreement between Llama-Guard-3-8B and human assessment of unsafe content, which to our knowledge is the largest human assessment of this type [Samvelyan et al.([n. d.]), Chao et al.(2024a)]. The outcomes of this work illustrate the current limits of model safety, and provide an accessible and simple methodology to reach state-of-the-art model safety with minimal resources.
@@ 2 Related work
Many contemporary open-source models incorporate DPO or its variants as a key component of their alignment pipelines [Intel([n. d.]), Zhu et al.(2024), Tunstall et al.(2023)]. While initial models often used relatively modest datasets, current state-of-the-art models, such as Llama 3 [Llama Team(2024)], Qwen 2.5 [Qwen et al.(2025)] and Tulu 3 [Lambert et al.(2024)], now use significantly larger preference datasets, often in the millions, for post-training alignment. However, beyond the efforts made by large organizations, an important question remains open: What is the minimal data requirement for an effective DPO-based safety alignment? While DPO has been shown to achieve optimal performance in preference alignment tasks when using 5,000 to 10,000 training samples [Saeidi et al.(2024)], it is uncertain whether this phenomenon translates to model safety, particularly in the presence of jailbreaking attacks, and if this threshold can be further reduced. Notice such findings would increase the accessibility of this alignment technique.
@@ 3.1 Egida Dataset
All gathered samples were manually labeled by the authors into 27 fine-grained topics in a multilabeling fashion (i.e., every instance can have several ones). A list of all fine-grained topics within Egida, together with their frequency can be found in Figure 1. Since there is a significant imbalance among fine-grained topics, and considering how some of these are too small for analysis, the authors recommend aggregating topics into a higher level of abstraction when using the dataset. In this paper, we propose and use one such categorization drawing inspiration from previous works performing similar analyses [Team(2024b), Samvelyan et al.([n. d.])]. The mapping between both is presented at the top of Table 2.
These 2,949 labeled instances are expanded using 18 different jailbreaking attacks, originating from Chen et al. [Chen et al.(2024)], Shen et al. [Shen et al.(2024)], DeepInception [Li et al.(2024)] and ReNeLLM [Ding et al.(2024b)]. Two additional attack styles are implemented using Qwen 72B Chat [Bai et al.(2023a)]: Past tense [Andriushchenko and Flammarion(2024)] and technical report writing [Samvelyan et al.([n. d.])]. For this latter source, model refusals are filtered and removed using rule-based mechanisms. As a result, the complete Egida is composed of 61,830 unsafe instances [^4] Also including the samples before adding any jailbreaking attack..
@@ 3.3.1 Human Preference and Llama-Guard-3-8B
To measure the reliability of Llama-Guard classifications, and confirm that its labels align with human judgment, a human evaluation is conducted, testing the agreement with the model. Five of the authors of this paper, 3 men and 2 women, participated as human evaluators, collectively reviewing a total of 1,000 questions selected from Egida. This evaluation improves previous work [Samvelyan et al.([n. d.])] by assessing a larger dataset and involving more annotators for increased reliability.

### v2

@@ 1 Introduction
A common concern across LLM domains and applications is jailbreaking [Wei et al.(2024a), Huang et al.(2023)]; the introduction of malicious prompts with the purpose of leading the model towards producing unsafe content to requests that, without the attack prompt, would not responded to. As of today, a wide variety of jailbreaking methods exist [Yi et al.(2024)], and more will appear as all it takes to find them is inference access to an LLM and imagination. Considering the challenges jailbreaking represents for LLM safety, this is introduced into the DPO experimentation as an fundamental aspect to study.
@@
The consistency of the above experiments is validated through an independent study on the agreement between Llama-Guard-3-8B and human assessment of unsafe content, which to our knowledge is the largest human assessment of this type [Samvelyan et al.(2024), Chao et al.(2024a)]. The outcomes of this work illustrate the current limits of model safety, and provide an accessible and simple methodology to reach state-of-the-art model safety with minimal resources.
@@ 2 Related work
Many contemporary open-source models incorporate DPO or its variants as a key component of their alignment pipelines [Intel(2024), Zhu et al.(2024), Tunstall et al.(2023)]. While initial models often used relatively modest datasets, current state-of-the-art models, such as Llama 3 [Llama Team(2024)], Qwen 2.5 [Qwen et al.(2025)] and Tulu 3 [Lambert et al.(2024)], now use significantly larger preference datasets, often in the millions, for post-training alignment. However, beyond the efforts made by large organizations, an important question remains open: What is the minimal data requirement for an effective DPO-based safety alignment? While DPO has been shown to achieve optimal performance in preference alignment tasks when using 5,000 to 10,000 training samples [Saeidi et al.(2024)], it is uncertain whether this phenomenon translates to model safety, particularly in the presence of jailbreaking attacks, and if this threshold can be further reduced. Notice such findings would increase the accessibility of this alignment technique.
@@ 3.1 Egida Dataset
All gathered samples were manually labeled by the authors into 27 fine-grained topics in a multilabeling fashion (i.e., every instance can have several ones). A list of all fine-grained topics within Egida, together with their frequency can be found in Figure 1. Since there is a significant imbalance among fine-grained topics, and considering how some of these are too small for analysis, the authors recommend aggregating topics into a higher level of abstraction when using the dataset. In this paper, we propose and use one such categorization drawing inspiration from previous works performing similar analyses [Team(2024b), Samvelyan et al.(2024)]. The mapping between both is presented at the top of Table 2.
These 2,949 labeled instances are expanded using 18 different jailbreaking attacks, originating from Chen et al. [Chen et al.(2024)], Shen et al. [Shen et al.(2024)], DeepInception [Li et al.(2024)] and ReNeLLM [Ding et al.(2024b)]. Two additional attack styles are implemented using Qwen 72B Chat [Bai et al.(2023a)]: Past tense [Andriushchenko and Flammarion(2024)] and technical report writing [Samvelyan et al.(2024)]. For this latter source, model refusals are filtered and removed using rule-based mechanisms. As a result, the complete Egida is composed of 61,830 unsafe instances [^4] Also including the samples before adding any jailbreaking attack..
@@ 3.3.1 Human Preference and Llama-Guard-3-8B
To measure the reliability of Llama-Guard classifications, and confirm that its labels align with human judgment, a human evaluation is conducted, testing the agreement with the model. Five of the authors of this paper, 3 men and 2 women, participated as human evaluators, collectively reviewing a total of 1,000 questions selected from Egida. This evaluation improves previous work [Samvelyan et al.(2024)] by assessing a larger dataset and involving more annotators for increased reliability.